THE TORSION-ROTATION SPECTRUM OF D_{2}O_{2}

J.-M. Flaud, Laboratoire de Physique Moléculaire et Atmosphériqe, Université Pierre et Marie Curie et CNRS, 4 place Jussieu, Tour 13, 75252 Paris Cedex, France; \underline{J. W. Johns} and Z. Lu, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario K1A 0R6, Canada.

The torsion-rotation spectrum of D_{2}O_{2} has been recorded in the region from 25 to 600 cm^{−1} at a resolution (1/MOPD) of 0.003 cm^{−1} using a modified Bomem Fourier transform spectrometer. The aim of the study is to improve the torsional potential of D_{2}O_{2} and to locate the weak oxygen-oxygen stretching fundamental which should lie amongst the torsional levels. So far, the pairs of torsional levels with n = 0, 1 and 2 and one of the levels with n = 3 have been analysed with J and K values up to 30 and 11 respectively. The current state of the analysis will be described.